In a bicycle race,lionel gives robert a 500meter advantage. Also, Lionel starts 15minutes after Rober. If lionel bikes at 17km/h and robert at 14km/h, how long will it take lionel after he starts biking to overtake robert?
pleasee help thanks:)

Answer by jorel1380(3719)

14 km/h x 1/4 hour=14/4 kilometers=3.5 kilometers
3.5 kilometers + .5 kilometers=4 kilometers
17n=14n+4
3n=4
n=4/3 hour before lionel overtakes robert..